# Carleman Estimates and Controllability of Forward Stochastic Parabolic Equations with General Dynamic Boundary Conditions

### v1

@@ 3 Observability Problems
We first have that
$\widetilde{A}$ $=\Phi-\ell_{t}$
$=\frac{2}{\beta_{0}}\|\mathcal{A}\|_{\infty}(\mathcal{A}\nu\cdot\nu)|\nabla l|-\ell_{t}$
$=\frac{2}{\beta_{0}}\|\mathcal{A}\|_{\infty}\lambda\mu\varphi(\mathcal{A}\nu\cdot\nu)|\nabla\psi|+\lambda T\varphi^{2}O(e^{2\mu|\psi|_{\infty}}).$
Notice that $\nabla\psi=\partial_{\nu}\psi\,\nu$ on $\Gamma$, then it follows that
$\widetilde{A}=\frac{2}{\beta_{0}}\|\mathcal{A}\|_{\infty}\lambda\mu\varphi(\mathcal{A}\nu\cdot\nu)|\partial_{\nu}\psi|+\lambda T\varphi^{2}O(e^{2\mu|\psi|_{\infty}}).$
Since $(\mathcal{A}\nu\cdot\nu)\partial_{\nu}\psi=\partial_{\nu}^{\mathcal{A}}\psi$ and that $\partial_{\nu}\psi<0$ on $\Gamma$, we conclude that
$\widetilde{A}=-\frac{2}{\beta_{0}}\|\mathcal{A}\|_{\infty}\lambda\mu\varphi\partial_{\nu}^{\mathcal{A}}\psi+\lambda T\varphi^{2}O(e^{2\mu|\psi|_{\infty}}).$
On the other hand, we have that
$\widetilde{B}$ $=-2\widetilde{A}\Phi-\widetilde{A}_{t}$
$=-2\left[-\frac{2}{\beta_{0}}\|\mathcal{A}\|_{\infty}\lambda\mu\varphi\partial_{\nu}^{\mathcal{A}}\psi+\lambda T\varphi^{2}O(e^{2\mu|\psi|_{\infty}})\right]\left[-\frac{2}{\beta_{0}}\|\mathcal{A}\|_{\infty}\lambda\mu\varphi\partial_{\nu}^{\mathcal{A}}\psi\right]$
$\quad\;+\frac{2}{\beta_{0}}\|\mathcal{A}\|_{\infty}\lambda\mu\varphi_{t}\partial_{\nu}^{\mathcal{A}}\psi+\lambda T\varphi\varphi_{t}O(e^{2\mu|\psi|_{\infty}})$
$=-\frac{8}{\beta_{0}^{2}}\|\mathcal{A}\|^{2}_{\infty}\lambda^{2}\mu^{2}\varphi^{2}|\partial_{\nu}^{\mathcal{A}}\psi|^{2}+\lambda^{2}T\varphi^{3}O(\mu e^{2\mu|\psi|_{\infty}})+\lambda\varphi^{2}O(\mu)+\lambda T^{2}\varphi^{3}O(e^{4\mu|\psi|_{\infty}}).$
Hence, we deduce that
$\widetilde{B}=-\frac{8}{\beta_{0}^{2}}\|\mathcal{A}\|^{2}_{\infty}\lambda^{2}\mu^{2}\varphi^{2}|\partial_{\nu}^{\mathcal{A}}\psi|^{2}+\lambda^{2}T\varphi^{3}O(\mu e^{2\mu|\psi|_{\infty}})+\lambda T^{2}\varphi^{3}O(e^{4\mu|\psi|_{\infty}}).$
We are now in position to prove Lemma 2.1.
Proof of Lemma 2.1.
The proof is divided into three steps.
@@
$\begin{aligned} &2\mathbb{E}\iint_{Q}\sum_{j,k=1}^{N}c^{jk}h_{x_{j}}h_{x_{k}}dxdt+\mathbb{E}\iint_{Q}Bh^{2}dxdt\\ &\leq\mathbb{E}\iint_{Q}\theta^{2}F_{1}^{2}dxdt-2\mathbb{E}\iint_{Q}\nabla\cdot(\mathcal{A}\nabla h\,dh)dx+\mathbb{E}\iint_{Q}\theta^{2}AZ^{2}dxdt\\ &\quad\;+2\mathbb{E}\iint_{Q}\sum_{j,k=1}^{N}\Bigg[\sum_{j^{\prime},k^{\prime}=1}^{N}\big(2a^{jk}a^{j^{\prime}k^{\prime}}\ell_{x_{j^{\prime}}}h_{x_{j}}h_{x_{k^{\prime}}}-a^{jk}a^{j^{\prime}k^{\prime}}\ell_{x_{j}}h_{x_{j^{\prime}}}h_{x_{k^{\prime}}}\big)\\ &\qquad\qquad\qquad\quad\;\;\;-\Psi a^{jk}h_{x_{j}}h+a^{jk}\Big(A\ell_{x_{j}}+\frac{\Psi_{x_{j}}}{2}\Big)h^{2}\Bigg]_{x_{k}}dxdt.\end{aligned}$ (A.11)
Integrating (A.3) on $\Sigma$, taking expectation on both sides and using the fact that $\theta(0,\cdot)=\theta(T,\cdot)=0$ and recalling (1.1), we conclude that
$\begin{aligned} &\,\mathbb{E}\iint_{\Sigma}\widetilde{B}h_{\Gamma}^{2}d\sigma dt+2\beta_{0}\mathbb{E}\iint_{\Sigma}\Phi|\nabla_{\Gamma}h_{\Gamma}|^{2}d\sigma dt\\ &\leq\mathbb{E}\iint_{\Sigma}\theta^{2}F_{2}^{2}d\sigma dt+\mathbb{E}\iint_{\Sigma}\theta^{2}\widetilde{A}\widehat{Z}^{2}d\sigma dt-2\mathbb{E}\iint_{\Sigma}h_{\Gamma}\mathcal{A}_{\Gamma}\nabla_{\Gamma}h_{\Gamma}\cdot\nabla_{\Gamma}\Phi d\sigma dt\\ &\quad-\mathbb{E}\iint_{\Sigma}\theta^{2}(\mathcal{A}_{\Gamma})_{t}\nabla_{\Gamma}z_{\Gamma}\cdot\nabla_{\Gamma}z_{\Gamma}d\sigma dt+2\mathbb{E}\iint_{\Sigma}\theta^{2}\partial_{\nu}^{\mathcal{A}}z\nabla_{\Gamma}\cdot(\mathcal{A}_{\Gamma}\nabla_{\Gamma}z_{\Gamma})d\sigma dt\\ &\quad+2\mathbb{E}\iint_{\Sigma}\theta^{2}\widetilde{A}z_{\Gamma}\partial_{\nu}^{\mathcal{A}}zd\sigma dt.\end{aligned}$ (A.12)
Combining (A.11) and (A.12), we get that
@@
$\begin{aligned} &2\mathbb{E}\iint_{Q}\sum_{j,k=1}^{N}c^{jk}h_{x_{j}}h_{x_{k}}dxdt+\mathbb{E}\iint_{Q}Bh^{2}dxdt+\textbf{I}_{1}+\textbf{I}_{2}\\ &\leq\mathbb{E}\iint_{Q}\theta^{2}F_{1}^{2}dxdt+\mathbb{E}\iint_{\Sigma}\theta^{2}F_{2}^{2}d\sigma dt+\mathbb{E}\iint_{Q}\theta^{2}AZ^{2}dxdt+\sum_{i=3}^{9}\textbf{I}_{i}.\end{aligned}$ (A.13)
@@
From Lemma A.2, we first have that
$\begin{aligned} \textbf{I}_{1}=\mathbb{E}\iint_{\Sigma}\widetilde{B}h_{\Gamma}^{2}d\sigma dt=\mathbb{E}\iint_{\Sigma}\theta^{2}\bigg[&-\frac{8}{\beta_{0}^{2}}\|\mathcal{A}\|^{2}_{\infty}\lambda^{2}\mu^{2}\varphi^{2}|\partial_{\nu}^{\mathcal{A}}\psi|^{2}+\lambda^{2}T\varphi^{3}O(\mu e^{2\mu|\psi|_{\infty}})\\ &+\lambda T^{2}\varphi^{3}O(e^{4\mu|\psi|_{\infty}})\bigg]z_{\Gamma}^{2}d\sigma dt,\end{aligned}$
@@
$\begin{aligned} \textbf{I}_{1}\geq&-C\lambda^{2}\mu^{2}\mathbb{E}\iint_{\Sigma}\theta^{2}\varphi^{2}z_{\Gamma}^{2}d\sigma dt-C\lambda^{2}\mu Te^{2\mu|\psi|_{\infty}}\mathbb{E}\iint_{\Sigma}\theta^{2}\varphi^{3}z_{\Gamma}^{2}d\sigma dt\\ &-C\lambda T^{2}e^{4\mu|\psi|_{\infty}}\mathbb{E}\iint_{\Sigma}\theta^{2}\varphi^{3}z_{\Gamma}^{2}d\sigma dt.\end{aligned}$ (A.14)
@@
$\textbf{I}_{2}=2\beta_{0}\mathbb{E}\displaystyle\iint_{\Sigma}\Phi|\nabla_{\Gamma}v_{\Gamma}|^{2}d\sigma dt=4\lambda\mu\|\mathcal{A}\|_{\infty}\mathbb{E}\displaystyle\iint_{\Sigma}\theta^{2}\varphi|\partial_{\nu}^{\mathcal{A}}\psi||\nabla_{\Gamma}z_{\Gamma}|^{2}d\sigma dt,$
@@
$\textbf{I}_{2}=-4\lambda\mu\|\mathcal{A}\|_{\infty}\mathbb{E}\displaystyle\iint_{\Sigma}\theta^{2}\varphi\partial_{\nu}^{\mathcal{A}}\psi|\nabla_{\Gamma}z_{\Gamma}|^{2}d\sigma dt.$ (A.15)
@@
$\begin{aligned} \textbf{I}_{3}^{1}&=2\lambda^{3}\mu^{3}\mathbb{E}\iint_{\Sigma}\theta^{2}\varphi^{3}\partial_{\nu}\psi|\partial_{\nu}^{\mathcal{A}}\psi|^{2}z_{\Gamma}^{2}d\sigma dt+4\lambda^{2}\mu^{2}\mathbb{E}\iint_{\Sigma}\theta^{2}\varphi^{2}\partial_{\nu}\psi\partial_{\nu}^{\mathcal{A}}\psi z_{\Gamma}\partial_{\nu}^{\mathcal{A}}zd\sigma dt\\ &\quad+4\lambda\mu\mathbb{E}\iint_{\Sigma}\theta^{2}\varphi\partial_{\nu}\psi|\partial_{\nu}^{\mathcal{A}}z|^{2}d\sigma dt-2\lambda\mu\mathbb{E}\iint_{\Sigma}\theta^{2}\varphi\partial_{\nu}\psi(\mathcal{A}\nabla z\cdot\nabla z)(\mathcal{A}\nu\cdot\nu)d\sigma dt.\end{aligned}$ (A.16)
Using the inequity $4ab\leq\frac{8}{3}a^{2}+\frac{3}{2}b^{2}$ for the second term in the right-hand side of (A.16), we obtain that
@@
$\begin{aligned} \textbf{I}_{3}^{1}&\leq-\frac{2}{3}\lambda^{3}\mu^{3}\mathbb{E}\iint_{\Sigma}\theta^{2}\varphi^{3}\partial_{\nu}\psi|\partial_{\nu}^{\mathcal{A}}\psi|^{2}z_{\Gamma}^{2}d\sigma dt+\frac{1}{2}\lambda\mu\mathbb{E}\iint_{\Sigma}\theta^{2}\varphi\partial_{\nu}\psi|\partial_{\nu}^{\mathcal{A}}z|^{2}d\sigma dt\\ &\quad+2\lambda\mu\mathbb{E}\iint_{\Sigma}\theta^{2}\varphi\partial_{\nu}\psi|\partial_{\nu}^{\mathcal{A}}z|^{2}d\sigma dt-2\lambda\mu\mathbb{E}\iint_{\Sigma}\theta^{2}\varphi\partial_{\nu}\psi(\mathcal{A}\nabla z\cdot\nabla z)(\mathcal{A}\nu\cdot\nu)d\sigma dt.\end{aligned}$ (A.17)
To estimate the last two terms on the right-hand side of (A.17), we use the identity (see [1, Lemma 2.3])
@@
$\begin{aligned} &2\lambda\mu\mathbb{E}\iint_{\Sigma}\theta^{2}\varphi\partial_{\nu}\psi|\partial_{\nu}^{\mathcal{A}}z|^{2}d\sigma dt-2\lambda\mu\mathbb{E}\iint_{\Sigma}\theta^{2}\varphi\partial_{\nu}\psi(\mathcal{A}\nabla z\cdot\nabla z)(\mathcal{A}\nu\cdot\nu)d\sigma dt\\ &\leq-2\lambda\mu\|\mathcal{A}\|_{\infty}\mathbb{E}\iint_{\Sigma}\theta^{2}\varphi\partial_{\nu}^{\mathcal{A}}\psi|\nabla_{\Gamma}z_{\Gamma}|^{2}d\sigma dt.\end{aligned}$ (A.18)
From (A.17) and (A.18), we deduce that
$\begin{aligned} \textbf{I}_{3}^{1}&\leq-\frac{2}{3}\lambda^{3}\mu^{3}\mathbb{E}\iint_{\Sigma}\theta^{2}\varphi^{3}\partial_{\nu}\psi|\partial_{\nu}^{\mathcal{A}}\psi|^{2}z_{\Gamma}^{2}d\sigma dt+\frac{1}{2}\lambda\mu\mathbb{E}\iint_{\Sigma}\theta^{2}\varphi\partial_{\nu}\psi|\partial_{\nu}^{\mathcal{A}}z|^{2}d\sigma dt\\ &\quad-2\lambda\mu\|\mathcal{A}\|_{\infty}\mathbb{E}\iint_{\Sigma}\theta^{2}\varphi\partial_{\nu}^{\mathcal{A}}\psi|\nabla_{\Gamma}z_{\Gamma}|^{2}d\sigma dt.\end{aligned}$ (A.19)
@@
$\begin{aligned} \textbf{I}_{3}^{2}\leq C\lambda\mu^{2}\mathbb{E}\iint_{\Sigma}\theta^{2}\varphi|z_{\Gamma}||\partial_{\nu}^{\mathcal{A}}z|d\sigma dt+C\lambda^{2}\mu^{2}\mathbb{E}\iint_{\Sigma}\theta^{2}\varphi^{2}z_{\Gamma}^{2}d\sigma dt.\end{aligned}$ (A.20)
Applying Young’s inequality for the first term on the right-hand side of (A.20), we have that
$\begin{aligned} \textbf{I}_{3}^{2}\leq C\mu\mathbb{E}\iint_{\Sigma}\theta^{2}\varphi|\partial_{\nu}^{\mathcal{A}}z|^{2}d\sigma dt+C\lambda^{2}\mu^{3}\mathbb{E}\iint_{\Sigma}\theta^{2}\varphi z_{\Gamma}^{2}d\sigma dt+C\lambda^{2}\mu^{2}\mathbb{E}\iint_{\Sigma}\theta^{2}\varphi^{2}z_{\Gamma}^{2}d\sigma dt.\end{aligned}$ (A.21)
@@
$\begin{aligned} \textbf{I}_{3}^{3}&=2\mathbb{E}\iint_{\Sigma}(A\lambda\mu\varphi\partial_{\nu}\psi\mathcal{A}\nu\cdot\nu+\frac{1}{2}\partial_{\nu}^{\mathcal{A}}\Psi)\theta^{2}z_{\Gamma}^{2}d\sigma dt\\ &=2\mathbb{E}\iint_{\Sigma}\left[\lambda^{3}\mu^{3}\theta^{2}\varphi^{3}z_{\Gamma}^{2}\partial_{\nu}\psi(\mathcal{A}\nu\cdot\nu)(\mathcal{A}\nabla\psi\cdot\nabla\psi)+\lambda^{2}\mu T\varphi^{3}O(e^{2\mu|\psi|_{\infty}})\right]d\sigma dt,\end{aligned}$
which implies for large enough $\lambda\geq CTe^{2\mu|\psi|_{\infty}}$,
$\begin{aligned} \textbf{I}_{3}^{3}\leq\frac{3}{2}\lambda^{3}\mu^{3}\mathbb{E}\iint_{\Sigma}\theta^{2}\varphi^{3}\partial_{\nu}\psi|\partial_{\nu}^{\mathcal{A}}\psi|^{2}z_{\Gamma}^{2}d\sigma dt.\end{aligned}$ (A.22)

### v2

@@ 3 Intermediate Carleman Estimate without Divergence Terms
We begin by recalling the definition $\widetilde{A}=\Phi-\ell_{t}$. By the choice of $\Phi$ and the properties of $\ell$, we have
$\widetilde{A}$ $=\frac{2}{\beta_{0}}M(\mathcal{A}\nu\cdot\nu)\,|\nabla\ell|-\ell_{t}$
$=\frac{2}{\beta_{0}}M\lambda\mu\varphi(\mathcal{A}\nu\cdot\nu)\,|\nabla\psi|+\lambda T\varphi^{2}\,O\!\left(e^{2\mu\|\psi\|_{\infty}}\right).$
Since $\nabla\psi=(\partial_{\nu}\psi)\nu$ on $\Gamma$, it follows that
$(\mathcal{A}\nu\cdot\nu)\,|\nabla\psi|=(\mathcal{A}\nu\cdot\nu)\,|\partial_{\nu}\psi|.$
Moreover, using the identity $\partial_{\nu}^{\mathcal{A}}\psi=(\mathcal{A}\nu\cdot\nu)\partial_{\nu}\psi$ and the fact that $\partial_{\nu}\psi<0$ on $\Gamma$, we obtain
$(\mathcal{A}\nu\cdot\nu)\,|\partial_{\nu}\psi|=-\partial_{\nu}^{\mathcal{A}}\psi.$
Consequently,
$\widetilde{A}=-\frac{2}{\beta_{0}}M\lambda\mu\varphi\,\partial_{\nu}^{\mathcal{A}}\psi+\lambda T\varphi^{2}\,O\!\left(e^{2\mu\|\psi\|_{\infty}}\right).$
We next estimate $\widetilde{B}$. By definition,
$\widetilde{B}=-2\widetilde{A}\Phi-\widetilde{A}_{t}.$
Substituting the above expression for $\widetilde{A}$ and recalling that $\Phi=-\frac{2}{\beta_{0}}M\lambda\mu\varphi\,\partial_{\nu}^{\mathcal{A}}\psi$, we compute
$\widetilde{B}$ $=-2\left(-\frac{2}{\beta_{0}}M\lambda\mu\varphi\,\partial_{\nu}^{\mathcal{A}}\psi+\lambda T\varphi^{2}\,O\!\left(e^{2\mu\|\psi\|_{\infty}}\right)\right)\left(-\frac{2}{\beta_{0}}M\lambda\mu\varphi\,\partial_{\nu}^{\mathcal{A}}\psi\right)$
$\quad-\partial_{t}\!\left(-\frac{2}{\beta_{0}}M\lambda\mu\varphi\,\partial_{\nu}^{\mathcal{A}}\psi+\lambda T\varphi^{2}\,O\!\left(e^{2\mu\|\psi\|_{\infty}}\right)\right).$
Using the bounds for $\varphi_{t}$ in (3.11), we deduce
$\widetilde{B}$ $=-\frac{8}{\beta_{0}^{2}}M^{2}\lambda^{2}\mu^{2}\varphi^{2}|\partial_{\nu}^{\mathcal{A}}\psi|^{2}+\lambda^{2}T\varphi^{3}\,O\!\left(\mu e^{2\mu\|\psi\|_{\infty}}\right)$
$\quad+\lambda T^{2}\varphi^{3}\,O\!\left(e^{4\mu\|\psi\|_{\infty}}\right),$
which completes the proof. ∎
As a consequence of the weighted identities and estimates established above, we obtain the following Carleman estimate for system (3.1).
Lemma 3.3.
[…]
$\begin{aligned} &\,\lambda^{3}\mu^{4}\mathbb{E}\iint_{Q}\theta^{2}\varphi^{3}|z|^{2}\,dxdt+\lambda^{3}\mu^{3}\mathbb{E}\iint_{\Sigma}\theta^{2}\varphi^{3}|z_{\Gamma}|^{2}\,d\sigma dt\\ &+\lambda\mu^{2}\mathbb{E}\iint_{Q}\theta^{2}\varphi|\nabla z|^{2}\,dxdt+\lambda\mu\mathbb{E}\iint_{\Sigma}\theta^{2}\varphi|\nabla_{\Gamma}z_{\Gamma}|^{2}\,d\sigma dt\\ &\leq C\bigg[\lambda^{3}\mu^{4}\mathbb{E}\int_{0}^{T}\int_{\mathcal{B}}\theta^{2}\varphi^{3}|z|^{2}\,dxdt+\mathbb{E}\iint_{Q}\theta^{2}|F_{1}|^{2}\,dxdt+\mathbb{E}\iint_{\Sigma}\theta^{2}|F_{2}|^{2}\,d\sigma dt\\ &\qquad\quad+\lambda^{2}\mu^{2}\mathbb{E}\iint_{Q}\theta^{2}\varphi^{2}|Z|^{2}\,dxdt+\lambda^{2}\mu\mathbb{E}\iint_{\Sigma}\theta^{2}\varphi^{2}|\widehat{Z}|^{2}\,d\sigma dt\bigg],\end{aligned}$ (3.12)
for large enough $\mu\geq C$ and $\lambda\geq C(e^{2\mu\|\psi\|_{\infty}}T+T^{2})$.
Proof.
For clarity, we divide the proof into three steps.
@@
$\begin{aligned} &2\mathbb{E}\iint_{Q}\sum_{j,k=1}^{N}c^{jk}h_{x_{j}}h_{x_{k}}dxdt+\mathbb{E}\iint_{Q}Bh^{2}dxdt\\ &\leq\mathbb{E}\iint_{Q}\theta^{2}F_{1}^{2}dxdt-2\mathbb{E}\iint_{Q}\nabla\cdot(\mathcal{A}\nabla h\,dh)dx+\mathbb{E}\iint_{Q}\theta^{2}AZ^{2}dxdt\\ &\quad\;+2\mathbb{E}\iint_{Q}\sum_{j,k=1}^{N}\bigg[\sum_{j^{\prime},k^{\prime}=1}^{N}\big(2a^{jk}a^{j^{\prime}k^{\prime}}\ell_{x_{j^{\prime}}}h_{x_{j}}h_{x_{k^{\prime}}}-a^{jk}a^{j^{\prime}k^{\prime}}\ell_{x_{j}}h_{x_{j^{\prime}}}h_{x_{k^{\prime}}}\big)\\ &\qquad\qquad\qquad\quad\;\;\;-\Psi a^{jk}h_{x_{j}}h+a^{jk}\Big(A\ell_{x_{j}}+\frac{\Psi_{x_{j}}}{2}\Big)h^{2}\bigg]_{x_{k}}dxdt.\end{aligned}$ (3.13)
Integrating (3.4) on $\Sigma$, taking expectation on both sides and using the fact that $\theta(0,\cdot)=\theta(T,\cdot)=0$ and recalling (1.1), we conclude that
$\begin{aligned} &\,\mathbb{E}\iint_{\Sigma}\widetilde{B}h_{\Gamma}^{2}d\sigma dt+2\beta_{0}\mathbb{E}\iint_{\Sigma}\Phi|\nabla_{\Gamma}h_{\Gamma}|^{2}d\sigma dt\\ &\leq\mathbb{E}\iint_{\Sigma}\theta^{2}F_{2}^{2}d\sigma dt+\mathbb{E}\iint_{\Sigma}\theta^{2}\widetilde{A}\widehat{Z}^{2}d\sigma dt-2\mathbb{E}\iint_{\Sigma}h_{\Gamma}\mathcal{A}_{\Gamma}\nabla_{\Gamma}h_{\Gamma}\cdot\nabla_{\Gamma}\Phi d\sigma dt\\ &\quad-\mathbb{E}\iint_{\Sigma}\theta^{2}(\mathcal{A}_{\Gamma})_{t}\nabla_{\Gamma}z_{\Gamma}\cdot\nabla_{\Gamma}z_{\Gamma}d\sigma dt+2\mathbb{E}\iint_{\Sigma}\theta^{2}\partial_{\nu}^{\mathcal{A}}z\nabla_{\Gamma}\cdot(\mathcal{A}_{\Gamma}\nabla_{\Gamma}z_{\Gamma})d\sigma dt\\ &\quad+2\mathbb{E}\iint_{\Sigma}\theta^{2}\widetilde{A}z_{\Gamma}\partial_{\nu}^{\mathcal{A}}zd\sigma dt.\end{aligned}$ (3.14)
Combining (3.13) and (3.14), we get that
@@
$\begin{aligned} &2\mathbb{E}\iint_{Q}\sum_{j,k=1}^{N}c^{jk}h_{x_{j}}h_{x_{k}}dxdt+\mathbb{E}\iint_{Q}Bh^{2}dxdt+\textbf{I}_{1}+\textbf{I}_{2}\\ &\leq\mathbb{E}\iint_{Q}\theta^{2}F_{1}^{2}dxdt+\mathbb{E}\iint_{\Sigma}\theta^{2}F_{2}^{2}d\sigma dt+\mathbb{E}\iint_{Q}\theta^{2}AZ^{2}dxdt+\sum_{i=3}^{9}\textbf{I}_{i}.\end{aligned}$ (3.15)
@@
From Lemma 3.2, we first have that
$\begin{aligned} \textbf{I}_{1}=\mathbb{E}\iint_{\Sigma}\widetilde{B}h_{\Gamma}^{2}d\sigma dt=\mathbb{E}\iint_{\Sigma}\theta^{2}\bigg[&-\frac{8}{\beta_{0}^{2}}M^{2}\lambda^{2}\mu^{2}\varphi^{2}|\partial_{\nu}^{\mathcal{A}}\psi|^{2}+\lambda^{2}T\varphi^{3}O(\mu e^{2\mu\|\psi\|_{\infty}})\\ &+\lambda T^{2}\varphi^{3}O(e^{4\mu\|\psi\|_{\infty}})\bigg]z_{\Gamma}^{2}d\sigma dt,\end{aligned}$
@@
$\begin{aligned} \textbf{I}_{1}\geq&-C\lambda^{2}\mu^{2}\mathbb{E}\iint_{\Sigma}\theta^{2}\varphi^{2}z_{\Gamma}^{2}d\sigma dt-C\lambda^{2}\mu Te^{2\mu\|\psi\|_{\infty}}\mathbb{E}\iint_{\Sigma}\theta^{2}\varphi^{3}z_{\Gamma}^{2}d\sigma dt\\ &-C\lambda T^{2}e^{4\mu\|\psi\|_{\infty}}\mathbb{E}\iint_{\Sigma}\theta^{2}\varphi^{3}z_{\Gamma}^{2}d\sigma dt.\end{aligned}$ (3.16)
@@
$\textbf{I}_{2}=2\beta_{0}\mathbb{E}\displaystyle\iint_{\Sigma}\Phi|\nabla_{\Gamma}v_{\Gamma}|^{2}d\sigma dt=4\lambda\mu M\mathbb{E}\displaystyle\iint_{\Sigma}\theta^{2}\varphi|\partial_{\nu}^{\mathcal{A}}\psi||\nabla_{\Gamma}z_{\Gamma}|^{2}d\sigma dt,$
@@
$\textbf{I}_{2}=-4\lambda\mu M\mathbb{E}\displaystyle\iint_{\Sigma}\theta^{2}\varphi\partial_{\nu}^{\mathcal{A}}\psi|\nabla_{\Gamma}z_{\Gamma}|^{2}d\sigma dt.$ (3.17)
@@
$\begin{aligned} \textbf{I}_{3}^{1}&=2\lambda^{3}\mu^{3}\mathbb{E}\iint_{\Sigma}\theta^{2}\varphi^{3}\partial_{\nu}\psi|\partial_{\nu}^{\mathcal{A}}\psi|^{2}z_{\Gamma}^{2}d\sigma dt+4\lambda^{2}\mu^{2}\mathbb{E}\iint_{\Sigma}\theta^{2}\varphi^{2}\partial_{\nu}\psi\partial_{\nu}^{\mathcal{A}}\psi z_{\Gamma}\partial_{\nu}^{\mathcal{A}}zd\sigma dt\\ &\quad+4\lambda\mu\mathbb{E}\iint_{\Sigma}\theta^{2}\varphi\partial_{\nu}\psi|\partial_{\nu}^{\mathcal{A}}z|^{2}d\sigma dt-2\lambda\mu\mathbb{E}\iint_{\Sigma}\theta^{2}\varphi\partial_{\nu}\psi(\mathcal{A}\nabla z\cdot\nabla z)(\mathcal{A}\nu\cdot\nu)d\sigma dt.\end{aligned}$ (3.18)
Using the inequity $4ab\leq\frac{8}{3}a^{2}+\frac{3}{2}b^{2}$ for the second term in the right-hand side of (3.18), we obtain that
@@
$\begin{aligned} \textbf{I}_{3}^{1}&\leq-\frac{2}{3}\lambda^{3}\mu^{3}\mathbb{E}\iint_{\Sigma}\theta^{2}\varphi^{3}\partial_{\nu}\psi|\partial_{\nu}^{\mathcal{A}}\psi|^{2}z_{\Gamma}^{2}d\sigma dt+\frac{1}{2}\lambda\mu\mathbb{E}\iint_{\Sigma}\theta^{2}\varphi\partial_{\nu}\psi|\partial_{\nu}^{\mathcal{A}}z|^{2}d\sigma dt\\ &\quad+2\lambda\mu\mathbb{E}\iint_{\Sigma}\theta^{2}\varphi\partial_{\nu}\psi|\partial_{\nu}^{\mathcal{A}}z|^{2}d\sigma dt-2\lambda\mu\mathbb{E}\iint_{\Sigma}\theta^{2}\varphi\partial_{\nu}\psi(\mathcal{A}\nabla z\cdot\nabla z)(\mathcal{A}\nu\cdot\nu)d\sigma dt.\end{aligned}$ (3.19)
To estimate the last two terms on the right-hand side of (3.19), we use the identity (see [1, Lemma 2.3])
@@
$\begin{aligned} &2\lambda\mu\mathbb{E}\iint_{\Sigma}\theta^{2}\varphi\partial_{\nu}\psi|\partial_{\nu}^{\mathcal{A}}z|^{2}d\sigma dt-2\lambda\mu\mathbb{E}\iint_{\Sigma}\theta^{2}\varphi\partial_{\nu}\psi(\mathcal{A}\nabla z\cdot\nabla z)(\mathcal{A}\nu\cdot\nu)d\sigma dt\\ &\leq-2\lambda\mu M\mathbb{E}\iint_{\Sigma}\theta^{2}\varphi\partial_{\nu}^{\mathcal{A}}\psi|\nabla_{\Gamma}z_{\Gamma}|^{2}d\sigma dt.\end{aligned}$ (3.20)
From (3.19) and (3.20), we deduce that
$\begin{aligned} \textbf{I}_{3}^{1}&\leq-\frac{2}{3}\lambda^{3}\mu^{3}\mathbb{E}\iint_{\Sigma}\theta^{2}\varphi^{3}\partial_{\nu}\psi|\partial_{\nu}^{\mathcal{A}}\psi|^{2}z_{\Gamma}^{2}d\sigma dt+\frac{1}{2}\lambda\mu\mathbb{E}\iint_{\Sigma}\theta^{2}\varphi\partial_{\nu}\psi|\partial_{\nu}^{\mathcal{A}}z|^{2}d\sigma dt\\ &\quad-2\lambda\mu M\mathbb{E}\iint_{\Sigma}\theta^{2}\varphi\partial_{\nu}^{\mathcal{A}}\psi|\nabla_{\Gamma}z_{\Gamma}|^{2}d\sigma dt.\end{aligned}$ (3.21)
@@
$\begin{aligned} \textbf{I}_{3}^{2}\leq C\lambda\mu^{2}\mathbb{E}\iint_{\Sigma}\theta^{2}\varphi|z_{\Gamma}||\partial_{\nu}^{\mathcal{A}}z|d\sigma dt+C\lambda^{2}\mu^{2}\mathbb{E}\iint_{\Sigma}\theta^{2}\varphi^{2}z_{\Gamma}^{2}d\sigma dt.\end{aligned}$ (3.22)
Applying Young’s inequality for the first term on the right-hand side of (3.22), we have that
$\begin{aligned} \textbf{I}_{3}^{2}\leq C\mu\mathbb{E}\iint_{\Sigma}\theta^{2}\varphi|\partial_{\nu}^{\mathcal{A}}z|^{2}d\sigma dt+C\lambda^{2}\mu^{3}\mathbb{E}\iint_{\Sigma}\theta^{2}\varphi z_{\Gamma}^{2}d\sigma dt+C\lambda^{2}\mu^{2}\mathbb{E}\iint_{\Sigma}\theta^{2}\varphi^{2}z_{\Gamma}^{2}d\sigma dt.\end{aligned}$ (3.23)
@@
$\begin{aligned} \textbf{I}_{3}^{3}&=2\mathbb{E}\iint_{\Sigma}(A\lambda\mu\varphi\partial_{\nu}\psi\mathcal{A}\nu\cdot\nu+\frac{1}{2}\partial_{\nu}^{\mathcal{A}}\Psi)\theta^{2}z_{\Gamma}^{2}d\sigma dt\\ &=2\mathbb{E}\iint_{\Sigma}\left[\lambda^{3}\mu^{3}\theta^{2}\varphi^{3}z_{\Gamma}^{2}\partial_{\nu}\psi(\mathcal{A}\nu\cdot\nu)(\mathcal{A}\nabla\psi\cdot\nabla\psi)+\lambda^{2}\mu T\varphi^{3}O(e^{2\mu\|\psi\|_{\infty}})\right]d\sigma dt,\end{aligned}$
which implies for large enough $\lambda\geq CTe^{2\mu\|\psi\|_{\infty}}$,
$\begin{aligned} \textbf{I}_{3}^{3}\leq\frac{3}{2}\lambda^{3}\mu^{3}\mathbb{E}\iint_{\Sigma}\theta^{2}\varphi^{3}\partial_{\nu}\psi|\partial_{\nu}^{\mathcal{A}}\psi|^{2}z_{\Gamma}^{2}d\sigma dt.\end{aligned}$ (3.24)